OpenStudy (anonymous):

Indefinite Integral using partial fraction decomposition method: Here's what I'm supposed to integrate: (2x+1)/(2x^2+4x) and my work: (2x+1)/(2x^2+4x) (2x+1)/(2x(x+2)) (2x+1)/(2x(x+2))= (A/2x)+(B/(x+2)) (2x+1)/(2x(x+2))= ((A)(x+2)+(B(2x))/(2x(x+2) (2x+1) = 2Ax +B(x+2) this is where I'm stuck, as you can see it's the algebraic steps that I'm having trouble with more than anything. Thanks for the help!

OpenStudy (anonymous):

instead of getting the A's and B's separated you need to get the x and the constant separated and then you just compare the coefficients and you will have 2 simultaneous equations with 2 unknowns

OpenStudy (anonymous):

(2x+1)= x(2A+B) + (2B) so 2A+B=2 and 2B=1

OpenStudy (anonymous):

Thanks a lot for the prompt response! I feel silly, as it's been a while since I've done any math using a system of equations (if that's the right word to use) but can you show me how (2x+1)=x(2A+B)+2B reduces down to 2A+B=2 and 2B=1 I also got B = 1/2 before using a guess and check method, and A = 3/4, which creates an easy-ish equation to integrate but it still wasn't the antiderivative of the opening problem so I felt I went wrong on this step. so if you could elaborate on what you wrote that would be a huge help. Thanks!!

OpenStudy (anonymous):

lemme do it nicely for you quick one second

OpenStudy (anonymous):

ok thank you adbermie. Take your time

OpenStudy (anonymous):

\[\frac{ 2x+1 }{ 2x^2+4x }=\frac{ 2x +1}{ 2x(x+2) }\] \[\frac{ 2x+1 }{ 2x(x+2 )}=\frac{ A }{ 2x }+\frac{ B }{ x+2 }\] multiply by the denominator of the left hand side \[2x+1=A(x+2)+B(2x)=(A+2B)x+2A\] and now you just need to make sure the coefficient of x and the constant are the same on both sides. it seems you just made a strange arithmetic error in your working so you end up with \[2A+B=2 \] and \[2A=1\] ie. \[A=\frac{ 1 }{ 2 }\] and thus \[B=1\]

OpenStudy (anonymous):

sorry haha i made a stupid mistake

OpenStudy (anonymous):

should be A+2B=2

OpenStudy (anonymous):

so B=3/4

OpenStudy (anonymous):

what did the answers give you as the antiderivative?

OpenStudy (anonymous):

hey well i appreciate you writing that all out for me adbermie, i understand how to get A & B now. the antiderivative i'm getting is 1/2ln|x+2| + 3/8ln|x|+C, but that doesnt derive to what we start with. once again thanks a lot for your help!

OpenStudy (anonymous):

sorry for responding a little late my internet connection was spotty

hartnn (hartnn):

did u get A=1/4, B=3/4 ?

OpenStudy (anonymous):

nope, I got A =1/2, B=3/4

hartnn (hartnn):

sorry A=1/2

hartnn (hartnn):

yeah thats correct

hartnn (hartnn):

now just put it in (2x+1)/(2x(x+2))= (A/2x)+(B/(x+2))

OpenStudy (anonymous):

Try using integration by recognition. The derivative of the bottom term is almost on the top, which you can force with some alegbra

OpenStudy (anonymous):

hey thanks for the responses guys. hartnn- shouldnt A and B be swapped? so its A/x+2 + B/2x thanks for the suggestion ZhangA. I'm not seeing how that would work though off the bat..

hartnn (hartnn):

why swapping? NO (2x+1)/(2x(x+2))= (A/2x)+(B/(x+2)) = 1/4x + 3/4(x+2) and thats it, just integrate

OpenStudy (anonymous):

haha, I wasnt swapping for the sake of swapping. Above I ended up with the terms A/x+2 + B/x+2...but the step before that was vice versa, like you have. let me see tho

OpenStudy (anonymous):

alright cool! thanks hartnn. I'm getting the same answer as wolframalpha